# Adaptive Teaming in Multi-Drone Pursuit: Simulation, Training, and Deployment

### v1

@@ 3.2 Simulation and Deployment
Simulation. The simulation module offers a user-friendly and customizable framework through the environment configurator and interface, enabling seamless configuration of multi-drone pursuit scenarios. Fig. 2 (bottom left) shows three examples of custom environments with photos from our real deployed system.
The environment configurator enables users to define all aspects of the simulation by categorizing configuration parameters into three key sections: players, site, and task, as shown in Fig. 7 in Appendix D. This modular design ensures flexibility for a variety of experimental setups. The players section configures participants, including learners, unseen teammates, and evaders, specifying parameters such as their numbers, velocities, and the inclusion of an unseen drone zoo. The site section allows users to customize the physical environment, including map dimensions and obstacle layouts, enabling diverse experimental terrains. Finally, the task section defines the task-specific parameters that govern the rules and objectives of the pursuit scenario.
Environment interface provides a Gymnasium-based interface for reinforcement learning-based training and evaluation. This interface ensures seamless interaction between the training algorithms and the simulation, enabling efficient development and testing of adaptive teaming strategies. By integrating these detailed configurations, the simulation module supports diverse and customized experimental setups while maintaining efficiency for training. Moreover, it bridges the gap between simulation and real-world multi-drone applications, facilitating a robust pipeline for research and deployment.
Deployment. The AT-MDP framework provides a robust real-world deployment solution, integrating edge computing nodes such as the Nvidia Jetson AGX Orin, the OptiTrack motion capture system, and CrazyFlie drones. As shown on the right of Fig. 2, the deployment workflow combines all components into a unified decision-making and execution pipeline. The learners’ policies and unseen drone partners, sampled from the unseen drone pool in the training module, are deployed on the edge nodes, which serve as inference engines for these policies. In this setup, the Jetson AGX Orin edge computing node acts as the core computational unit, offering compact, energy-efficient, on-site processing capabilities. It enables real-time policy inference to support efficient drone decision-making during pursuit tasks. Currently deployed alongside CrazyFlie drones, its lightweight design holds potential for future integration directly onto drones, enabling fully autonomous and decentralized multi-drone systems.
Observations from the physical environment, including drone positions and movements, are captured by the OptiTrack motion capture system, which provides high-precision localization and tracking. The processed observations are sent to the edge nodes, which supply real-time data to the inference engine. The inference engine interprets the data and generates appropriate actions based on the policies deployed. These actions are then communicated to the drones for real-time execution. The drones provide continuous feedback to the system, facilitating dynamic adjustments and enhancing coordination among agents.
Table 2: Performance comparison across different difficulty levels for adaptive teaming without teammate modeling. The means and standard deviations are calculated over five different random seeds, with each seed undergoing 50 repeated runs.
ENV Metrics Unseen Zoo 1 Unseen Zoo 2 Unseen Zoo 3 Unseen Zoo 4
PBT MAPPO OPT PBT MAPPO OPT PBT MAPPO OPT PBT MAPPO OPT
4p2e3o SUC $\uparrow$ 65.20 74.00 77.67 62.00 70.80 86.00 54.80 65.20 77.20 64.80 74.80 82.80
$\pm$ 17.75 $\pm$ 4.69 $\pm$ 6.81 $\pm$ 12.41 $\pm$ 9.23 $\pm$ 3.46 $\pm$ 11.54 $\pm$ 7.69 $\pm$ 3.03 $\pm$ 15.01 $\pm$ 5.40 $\pm$ 1.79
COL $\downarrow$ 32.40 25.60 22.33 27.60 22.40 13.20 39.20 34.00 22.40 31.60 24.00 16.80
$\pm$ 15.19 $\pm$ 4.34 $\pm$ 6.81 $\pm$ 13.30 $\pm$ 12.60 $\pm$ 4.15 $\pm$ 12.30 $\pm$ 7.35 $\pm$ 3.58 $\pm$ 14.38 $\pm$ 4.90 $\pm$ 2.28
AST $\downarrow$ 313.48 303.21 259.17 435.65 386.32 370.95 380.28 376.14 314.97 380.85 328.96 306.35
$\pm$ 86.38 $\pm$ 51.91 $\pm$ 34.28 $\pm$ 113.16 $\pm$ 29.18 $\pm$ 31.36 $\pm$ 68.19 $\pm$ 34.65 $\pm$ 27.68 $\pm$ 85.25 $\pm$ 34.36 $\pm$ 25.86
REW $\uparrow$ 123.60 132.34 138.28 130.51 136.40 149.89 114.00 121.98 135.71 126.15 133.11 143.55
$\pm$ 22.68 $\pm$ 4.68 $\pm$ 7.95 $\pm$ 8.77 $\pm$ 8.50 $\pm$ 3.28 $\pm$ 10.13 $\pm$ 9.07 $\pm$ 3.46 $\pm$ 13.39 $\pm$ 4.68 $\pm$ 1.29
4p2e1o SUC $\uparrow$ 68.80 69.20 80.80 63.20 64.40 73.20 59.20 53.20 67.60 62.40 63.20 75.60
$\pm$ 5.02 $\pm$ 11.71 $\pm$ 7.56 $\pm$ 15.27 $\pm$ 12.20 $\pm$ 3.03 $\pm$ 8.07 $\pm$ 14.11 $\pm$ 5.18 $\pm$ 3.85 $\pm$ 14.46 $\pm$ 5.55
COL $\downarrow$ 30.80 30.40 19.20 31.20 28.80 23.20 36.80 43.20 32.00 34.80 35.60 23.60
$\pm$ 5.22 $\pm$ 12.03 $\pm$ 7.56 $\pm$ 12.85 $\pm$ 15.53 $\pm$ 1.79 $\pm$ 10.73 $\pm$ 15.59 $\pm$ 5.83 $\pm$ 5.40 $\pm$ 14.52 $\pm$ 6.54
AST $\downarrow$ 352.24 317.55 298.93 399.87 383.30 356.65 446.67 395.92 353.35 360.08 375.41 313.60
$\pm$ 30.08 $\pm$ 25.07 $\pm$ 29.19 $\pm$ 86.90 $\pm$ 75.08 $\pm$ 26.78 $\pm$ 72.36 $\pm$ 51.69 $\pm$ 35.29 $\pm$ 48.52 $\pm$ 22.18 $\pm$ 34.18
REW $\uparrow$ 122.05 126.73 138.24 125.85 127.30 134.26 112.66 109.06 120.83 118.73 120.00 131.50
$\pm$ 6.91 $\pm$ 9.92 $\pm$ 6.28 $\pm$ 13.38 $\pm$ 10.62 $\pm$ 3.04 $\pm$ 13.42 $\pm$ 11.94 $\pm$ 6.84 $\pm$ 5.91 $\pm$ 13.22 $\pm$ 5.76
4p2e5o SUC $\uparrow$ 40.40 47.00 50.00 62.40 72.00 75.20 67.60 65.50 74.00 54.80 58.50 67.60
$\pm$ 2.61 $\pm$ 10.39 $\pm$ 11.64 $\pm$ 3.29 $\pm$ 3.65 $\pm$ 6.72 $\pm$ 14.59 $\pm$ 3.42 $\pm$ 4.24 $\pm$ 2.28 $\pm$ 5.97 $\pm$ 6.39
COL $\downarrow$ 59.60 53.00 49.86 35.20 24.00 22.80 31.20 33.50 25.20 44.00 40.50 32.00
$\pm$ 2.61 $\pm$ 10.39 $\pm$ 11.62 $\pm$ 3.90 $\pm$ 5.89 $\pm$ 7.43 $\pm$ 14.74 $\pm$ 3.42 $\pm$ 3.35 $\pm$ 2.45 $\pm$ 5.74 $\pm$ 6.16
AST $\downarrow$ 333.96 313.53 331.98 349.48 419.21 396.66 287.18 348.23 281.45 294.41 340.94 313.79
$\pm$ 34.83 $\pm$ 51.81 $\pm$ 80.30 $\pm$ 17.64 $\pm$ 50.81 $\pm$ 21.48 $\pm$ 18.91 $\pm$ 55.20 $\pm$ 40.77 $\pm$ 38.79 $\pm$ 34.95 $\pm$ 26.78
REW $\uparrow$ 90.16 93.16 99.31 116.19 128.45 134.02 120.48 124.92 128.45 107.56 107.08 119.93
$\pm$ 7.47 $\pm$ 12.00 $\pm$ 17.70 $\pm$ 8.28 $\pm$ 4.27 $\pm$ 6.12 $\pm$ 19.11 $\pm$ 8.33 $\pm$ 3.79 $\pm$ 5.92 $\pm$ 5.44 $\pm$ 6.11
4p3e5o SUC $\uparrow$ 34.80 34.80 31.14 50.40 59.20 67.20 57.20 54.80 60.40 42.00 54.80 60.00
$\pm$ 5.22 $\pm$ 5.40 $\pm$ 17.20 $\pm$ 7.27 $\pm$ 5.40 $\pm$ 5.76 $\pm$ 6.57 $\pm$ 7.43 $\pm$ 5.18 $\pm$ 6.78 $\pm$ 9.44 $\pm$ 6.00
COL $\downarrow$ 62.80 64.80 67.57 46.80 36.00 28.80 40.40 38.00 36.40 55.60 41.60 38.40
$\pm$ 4.60 $\pm$ 5.02 $\pm$ 15.85 $\pm$ 7.56 $\pm$ 4.69 $\pm$ 6.42 $\pm$ 8.05 $\pm$ 5.83 $\pm$ 4.34 $\pm$ 4.77 $\pm$ 8.65 $\pm$ 5.18
AST $\downarrow$ 431.44 509.56 459.55 457.68 489.79 394.81 446.07 555.10 407.85 425.94 510.96 416.91
$\pm$ 67.14 $\pm$ 11.17 $\pm$ 91.71 $\pm$ 33.64 $\pm$ 51.33 $\pm$ 48.65 $\pm$ 38.97 $\pm$ 49.11 $\pm$ 63.06 $\pm$ 60.10 $\pm$ 38.53 $\pm$ 76.21
REW $\uparrow$ 141.98 131.29 116.57 168.82 187.30 193.78 182.04 187.32 196.07 146.74 172.06 174.59
$\pm$ 9.66 $\pm$ 15.98 $\pm$ 43.94 $\pm$ 10.17 $\pm$ 9.08 $\pm$ 4.05 $\pm$ 7.17 $\pm$ 8.97 $\pm$ 9.91 $\pm$ 11.93 $\pm$ 17.01 $\pm$ 18.68
@@ 4 Experiment
In this section, we evaluate the performance of baseline methods from the algorithm zoo and validate their feasibility in real-world multi-drone pursuit scenarios. The experiments are structured into two main parts: (1) adaptive teaming without teammate modeling, detailed in Section 4.1, and (2) adaptive teaming with teammate modeling, discussed in Section 4.2. Each subsection presents the experimental setup, key results, and in-depth analysis. Additionally, Section 4.3 provides a case study demonstrating real-world deployment, showcasing how adaptive learners coordinate with unseen drone partners to execute a multi-stage capture strategy. Further details on the real-world implementation are available in Appendix E.
We begin by introducing the shared experimental setups applicable to both groups of experiments.
Environments. To systematically evaluate the performance of adaptive teaming, we leverage the environment configurator to design four multi-drone pursuit environments with varying levels of difficulty, denoted as 4p2e3o, 4p2e1o, 4p2e5o, and 4p3e5o. Fig. 6 in the Appendix gives the screenshots of the four environments deployed in real world. The notation represents the number of pursuers (p), evaders (e), and obstacles (o) in each setting. Each environment is initialized with defined spawn areas, where the evaders spawn within a 3.2m wide and 0.6m high region, while the pursuers spawn in a similarly sized region. The obstacle layouts introduce additional complexity. Environment 4p2e3o include three distributed barriers with different shapes: two cubes in the left area and one cylinder in the right area. Although Environment 4p2e1o only includes a cuboid obstacle in the middle, the difficulty is slight higher than 4p2e3o for evaders have large space to escape. In general, the difficulity of the two environments is easy. Those environments with 5o feature densely packed obstacles that significantly constrain movement. In terms of difficulty, 4p2e3o is categorized as easy due to fewer obstacles and ample space for pursuit, 4p2e1o presents a moderate challenge with a single strategically placed obstacle, 4p2e5o is considered hard as it requires advanced maneuvering in a confined space, and 4p3e5o is the most difficult (superhard) due to an additional evader and dense obstacle layout, requiring highly coordinated teamwork for successful pursuit.
Evaluation Metrics. To evaluate the performance of the methods, we use the following metrics: 1. Success rate (SUC), which measures the proportion of tasks successfully completed. An episode is deemed successful if the pursuers capture both evaders, defined as reducing the distance between an evader and a pursuer to less than 0.2 metres; 2. Collision rate (COL), which tracks the frequency of collisions during task execution. A collision is recorded if the distance between any two pursuers is less than 0.2 metres or if the distance between a drone and an obstacle is less than 0.1 metres; 3. Average success timesteps (AST), which indicates the average number of steps taken to complete a task; 4. Average reward (REW), which reflects the overall efficiency and quality of the agent’s performance across episodes. These metrics collectively provide a comprehensive evaluation of the proposed method’s adaptability and effectiveness.
4.1 Adaptive teaming without teammate modeling
4.1.1 Evaluation Protocol
To ensure a rigorous and fair comparison of adaptive teaming, we standardize the evaluation process across all methods. Since AT-MDP is a newly introduced problem with no existing algorithms specifically designed for it, we establish baselines using PBT (Carroll et al., 2019), MAPPO (Yu et al., 2022), and our proposed OPT.
Unseen Drone Partners. Unseen drone partners are a critical component of the evaluation protocol. To thoroughly assess the proposed method, we evaluate its performance across four distinct unseen drone zoos, with drone policies selected from the unseen drone pool in the AT-MDP framework. Each zoo represents a unique combination of drone behaviors, designed to create diverse and challenging scenarios that test the adaptability and robustness of the proposed method. (1) Unseen Zoo 1 and (2) Unseen Zoo 3 include a Greedy drone policy and a VICSEK drone policy (Janosov et al., 2017a), respectively, providing additional diversity in drone behavior and coordination strategies. (3) Unseen Zoo 3 consists of two drone policies trained using a self-play PPO algorithm (Schulman et al., 2017; Carroll et al., 2019) with different random seeds. These policies are selected to reflect varying levels of coordination ability: one with strong coordination skills and another with relatively weaker performance. For example, in a 4p2e1o scenario, the success rates (SUC) of the two policies, when evaluated against themselves, are 70% and 54%, respectively. (4) Unseen Zoo 4 combines all policies from the other three zoos into a single zoo. At the beginning of each episode, unseen drone partners are randomly selected from this combined pool, introducing additional variability and unpredictability to the evaluation process.
4.1.2 Main Results
Performance Across Different Environments. As the complexity of the environments increases (from 4p2e3o to 4p3e5o), the success rate (SUC) generally decreases across all methods, highlighting the increasing difficulty in coordinating with unseen teammates. In simpler environments like 4p2e3o and 4p2e1o, OPT consistently outperforms other baselines, achieving higher success rates and lower collision rates. In more challenging settings (4p2e5o and 4p3e5o), performance gaps become more pronounced, with MAPPO and PBT struggling to generalise effectively.
Effectiveness of OPT in Adaptive Teaming. OPT achieves the highest success rates in most cases, particularly in Unseen Zoo 2 and Unseen Zoo 4, demonstrating its robustness in adapting to diverse teammate behaviors. It significantly reduces AST, meaning it captures evaders faster than other baselines, suggesting that its learning process leads to more efficient pursuit strategies. Moreover, OPT consistently achieves higher reward values, indicating its superior adaptability in unseen teaming scenarios.
Impact of Unseen Drone Zoos. The performance of all methods varies significantly across different unseen drone zoos, each introducing distinct teammate behaviors. Unseen zoo 1 (Greedy agents) and Unseen zoo 2 (VICSEK-based agents) exhibit more predictable behaviors, allowing for more effective coordination, with all methods achieving relatively higher SUC scores. Unseen zoo 3, composed entirely of PPO-based self-play policies with two different skill levels, presents the most challenging setting, where all baselines experience performance drops due to the increased adaptability and unpredictability of teammates. Unseen zoo 4, which includes a mix of all drone types with random selection, further increases variability, testing the generalization ability of each method. Notably, OPT remains the most stable approach, demonstrating robust adaptability across diverse teammate dynamics.
@@ 4.1.2 Main Results
4.2 Adaptive teaming with teammate modeling
We also conduct a group of experiments about adatptive teaming with teammates modeling. As shown Table. 3 presents the performance comparison between MAPPO and ATM across different difficulty levels. As the environment complexity increases (from 4p2e3o to 4p3e5o), both methods exhibit a decline in success rate (SUC) and an increase in average steps to capture (AST), highlighting the growing challenge of coordinating with teammates in more constrained and adversarial settings. ATM consistently outperforms MAPPO in success rate across all environments, demonstrating its improved adaptability in dynamic multi-agent interactions. Notably, ATM achieves a significantly higher success rate in 4p2e3o and 4p2e1o, suggesting its effectiveness in simpler settings. However, as the complexity increases (4p2e5o and 4p3e5o), the performance gap narrows, indicating the growing challenge of maintaining robust coordination under more difficult conditions. In terms of collision rate (COL), ATM generally results in lower or comparable values to MAPPO, suggesting that its teammate modeling mechanism helps mitigate unnecessary collisions. The reward metric (REW) further supports these findings, as ATM achieves higher values across all settings, reflecting its superior learning of adaptive strategies.
4.3 Case Study
This case study demonstrates how the ATM learners effectively coordinate with their unseen drone partners to execute a multi-stage capture strategy. Fig. 3 illustrates key frames from the scenario. In Frame 1, four pursuers initiate a collaborative approach, positioning themselves strategically to encircle all three evaders while maintaining an adaptive formation. In Frame 2, two pursuers successfully capture one of the evaders while the other two tighten their formation, preventing the remaining evaders from escaping. In Frames 3 & 4, the remaining two evaders are captured one by one as the pursuers continue refining their positioning and coordination, effectively closing all escape routes.
Table 3: Performance comparison across different difficulty levels. Metrics are reported as mean $\pm$ standard deviation.
Metrics 4p2e3o 4p2e1o 4p2e5o 4p3e5o
MAPPO ATM MAPPO ATM MAPPO ATM MAPPO ATM
SUC $\uparrow$ 72.80 77.20 71.60 72.00 52.40 72.40 51.20 62.00
($\pm$ 4.15) ($\pm$ 6.87) ($\pm$ 3.29) ($\pm$ 4.00) ($\pm$ 5.55) ($\pm$ 5.55) ($\pm$ 7.95) ($\pm$ 8.00)
COL $\downarrow$ 26.40 22.00 26.40 26.40 47.60 26.80 48.00 38.00
($\pm$ 4.56) ($\pm$ 7.62) ($\pm$ 3.58) ($\pm$ 6.07) ($\pm$ 5.55) ($\pm$ 6.10) ($\pm$ 7.48) ($\pm$ 8.00)
AST $\downarrow$ 297.46 277.55 370.15 326.03 289.01 283.59 408.28 372.33
($\pm$ 37.98) ($\pm$ 17.55) ($\pm$ 65.15) ($\pm$ 48.37) ($\pm$ 26.45) ($\pm$ 45.80) (±($\pm$( ± 45.57) ($\pm$ 17.39)
REW $\uparrow$ 133.03 136.84 128.79 130.08 104.79 128.45 164.19 180.54
($\pm$ 4.21) ($\pm$ 6.91) ($\pm$ 1.79) ($\pm$ 4.71) ($\pm$ 3.76) ($\pm$ 6.17) ($\pm$ 14.67) ($\pm$ 15.67)
Figure 3: Case Study: This example demonstrates the capture strategy executed by ATM learners and unseen drone partners from unseen zoo 3 in the superhard environment 4p3e5o. The red circles denote pursuers, and the black squares represent evaders. In this scenario, four pursuers collaboratively surround all three evader (1), two pursuers capture one of evaders while other two pursuers continuously tighten their formation (2), and rest of two evaders are then successfully captured one by one (3 & 4)

### v2

@@ 3.2 Simulation and Deployment
Simulation. The simulation module provides a highly customizable and intuitive framework through the environment configurator and the Gymnasium-based environment interface, enabling efficient setup and execution of multi-drone pursuit scenarios. The environment configurator organizes simulation parameters into three distinct categories for easy customization: players, site, and task (see Fig. 5 in Appendix A). The players category specifies the numbers, velocities, and characteristics of learners, unseen teammates, and evaders, and optionally incorporates an unseen drone zoo to simulate varied adaptive teaming conditions. The site category allows users to adjust the simulation environment’s physical properties, including map dimensions and obstacle configurations, enabling a wide range of scenario complexities. The task category defines rules and objectives unique to each pursuit scenario, determining conditions for success and interaction dynamics.
Leveraging the configurator, we systematically design four progressive multi-drone pursuit environments in the benchmark, named 4p2e3o, 4p2e1o, 4p2e5o, and 4p3e5o, indicating the number of pursuers (p), evaders (e), and obstacles (o). Screenshots of these real-world environments are provided in Fig. 4 in the Appendix. Evaders spawn within a specified region measuring 3.2m wide and 0.6m high, with pursuers spawning in a similar area. Obstacle configurations vary significantly across environments to introduce different complexity levels. Environment 4p2e3o, considered easy, contains three distributed obstacles (two cubes and one cylinder), providing ample pursuit space. Although environment 4p2e1o has only a single central obstacle, it poses a slightly higher difficulty due to evaders having greater freedom of movement. Environments featuring five obstacles (4p2e5o and 4p3e5o) are challenging, requiring sophisticated maneuvering due to densely packed obstacles restricting drone movements. Specifically, 4p2e5o is categorized as hard, while 4p3e5o is identified as the most challenging scenario (superhard), necessitating advanced coordination strategies for successful adaptive teaming.
In these environments, each agent’s observation includes: (1) the relative position and bearing to each evader within its perception range, (2) the distance and angle to the nearest obstacle, and (3) the relative position and bearing of nearby teammates. Observations are structured as normalized continuous vectors, with masking applied to represent occluded entities clearly. The action space is continuous, defined within the range $[-1,1]$, directly corresponding to angular steering adjustments that control the drone’s directional rotation during pursuit. The reward function incentivizes efficient pursuit behaviors through positive rewards for capturing evaders and additional shaped rewards proportional to the agents’ proximity improvement towards targets. Safety is enforced via proximity-based penalties for approaching too close to obstacles or teammates, effectively discouraging risky behaviors and promoting cooperative, safe navigation.
Deployment. As shown on the right side of Fig. 1, the AT-Drone benchmark supports real-world deployment within a 3.6m×$\times$ 5m area by seamlessly integrating edge computing nodes—such as the Nvidia Jetson Orin Nano and personal laptops—with Crazyflie drones. Specifically, we utilize the FZMotion system to perform real-time position tracking, transmitting positional data in point cloud format to Crazyswarm, where it is processed and fed into the decision-making policies. Policies for both the adaptive learners and unseen drone partners (sampled from the unseen drone pool) run on edge computing nodes that serve as inference engines. The Crazyswarm platform and the adaptive teaming policies are deployed separately across two edge devices: a Lenovo ThinkPad T590 laptop and a Jetson Orin Nano. These nodes handle the reception of drone position data from the motion capture system, execute the adaptive teaming algorithms, and transmit control commands to the Crazyflie drones via Crazyradio PA. Upon receiving control signals, the Crazyflie drones execute the maneuvers using their onboard Mellinger controller, ensuring accurate and responsive trajectory tracking. Overall, this real-world deployment setup allows us to directly evaluate the applicability of learned policies on physical drone systems, effectively bridging the gap between simulation and real-world application.
3.3 Training and Evaluation
Training. As illustrated in Fig. 1, the AT-Drone framework employs a distributed training architecture leveraging multiple parallel environments to efficiently scale the learning processes. To the best of our knowledge, this is the first study specifically targeting adaptive teaming strategies in multi-pursuer multi-evader drone pursuit scenarios, highlighting the novelty and importance of establishing a comprehensive algorithm zoo for rigorous benchmarking and broader community adoption.
To further boost zero-shot coordination capabilities, AT-Drone employs self-play and PBT strategies. Self-play enables drones to iteratively refine policies via competitive interactions against progressively updated versions of themselves, significantly improving their adaptability and coordination. Simultaneously, PBT [25] facilitates extensive exploration and efficient knowledge sharing among diverse model populations, thus enhancing generalization to dynamic pursuit tasks. Both self-play and PBT are implemented using Independent PPO (IPPO) [37], providing a robust, scalable training environment ideal for real-world multi-drone collaboration.
Recently, the HOLA-Drone method [24] introduces a hypergraphical-form game to model single-learner scenarios involving interactions with multiple unseen teammates. However, this method faces significant limitations when extended to more complex multi-learner scenarios common in multi-drone pursuit tasks. To overcome these challenges, we propose HOLA-Drone V2, an enhanced and generalized approach. Our key improvements include the construction of a preference hypergraph to explicitly identify and retain optimal teammate interactions, combined with a novel max-min preference oracle. This oracle systematically identifies challenging teammate subsets and iteratively optimizes drone strategies to robustly handle these scenarios. By dynamically adjusting the learner strategy set, our method significantly improves adaptability and coordination effectiveness in realistic multi-agent environments. Comprehensive algorithmic details, formal definitions, and implementations are provided in Appendix C.
In addition to zero-shot coordination methods, we introduce an ad-hoc teamwork algorithm, NAHT-D (NAHT for Drones), based on the recent NAHT framework [21]. Unlike the original NAHT method for discrete-action tasks (e.g., SMAC), NAHT-D is adapted for continuous-action drone scenarios, crucial for realistic maneuvering and pursuit tasks. NAHT-D efficiently models unseen drone teammates by integrating a specialized teammate-modeling network into MAPPO [38]. The teammate-modeling network employs an autoencoder, taking the past $k$ steps of observations and actions to reconstruct teammates’ current action distributions. For continuous action spaces, we use KL divergence as the reconstruction loss. The encoder’s output embedding captures teammate behaviors and is combined with the agent’s current observations as input to the policy network. For detailed algorithmic implementation, please refer to Appendix D.
Evaluation. To rigorously evaluate adaptive teaming strategies within multi-drone pursuit scenarios, we design a structured evaluation protocol consisting of multiple distinct unseen drone zoos and clearly defined evaluation metrics.
We construct a set of unseen drone zoos to introduce diverse and challenging teammate behaviors, spanning rule-based, bio-inspired, and learning-based methods. Specifically, these zoos include: (1) the Greedy Drone, employing a straightforward pursuit strategy focused on the nearest evader while dynamically avoiding collisions; (2) the VICSEK Drone, inspired by swarm behaviors, optimizing collective drone movement for pursuit and collision avoidance; and (3) the Self-Play Drones, trained using randomized IPPO-based self-play to generate diverse and unpredictable coordination behaviors. To ensure thorough evaluation, we define three distinct configurations of unseen drone partners: Unseen Zoo 1: Greedy drones only, highlighting direct pursuit behaviors. Unseen Zoo 2: Includes two IPPO self-play drone policies demonstrating different coordination skill levels—one highly coordinated (70% success) and one less coordinated (54% success)—introducing variability in teaming performance. Unseen Zoo 3: Combines all drones from the previous zoos, randomly selecting partners each episode to maximize behavioral diversity and unpredictability. Detailed descriptions and implementations of these drone behaviors are provided in Appendix B.
We adopt four quantitative metrics to systematically evaluate adaptive teaming performance: Success Rate (SUC): Percentage of episodes successfully completed, defined by capturing both evaders within 0.2 meters. Collision Rate (COL): Frequency of collisions between drones (threshold of 0.2 meters) or drones and obstacles (threshold of 0.1 meters), assessing operational safety. Average Success Timesteps (AST): Mean number of timesteps required for successful task completion, indicating pursuit efficiency. Average Reward (REW): Overall quality and efficiency of agent performance across episodes.
Figure 2: Success rate (SUC) across different difficulty levels for adaptive teaming without teammate modelling. Red dotted lines denote best-response baselines specifically trained on the given unseen teammate zoo.
ENV Metrics Unseen Zoo 1 Unseen Zoo 2 Unseen Zoo 3
SP PBT H-D w/o g H-D SP PBT H-D w/o g H-D SP PBT H-D w/o g H-D
[…]
$\pm$4.34 $\pm$15.19 $\pm$7.27 $\pm$6.81 $\pm$7.35 $\pm$12.10 $\pm$4.15 $\pm$3.58 $\pm$4.90 $\pm$14.38 $\pm$8.79 $\pm$2.28
[…]
$\pm$51.91 $\pm$86.38 $\pm$20.68 $\pm$34.28 $\pm$34.65 $\pm$68.19 $\pm$23.35 $\pm$27.68 $\pm$34.36 $\pm$85.25 $\pm$21.18 $\pm$25.86
[…]
$\pm$4.68 $\pm$22.68 $\pm$5.75 $\pm$7.95 $\pm$9.07 $\pm$10.13 $\pm$4.78 $\pm$3.46 $\pm$4.68 $\pm$13.39 $\pm$9.48 $\pm$1.29
[…]
$\pm$12.03 $\pm$5.22 $\pm$7.35 $\pm$7.56 $\pm$15.59 $\pm$10.73 $\pm$8.29 $\pm$5.83 $\pm$14.52 $\pm$5.40 $\pm$6.07 $\pm$6.54
[…]
$\pm$25.07 $\pm$30.08 $\pm$29.16 $\pm$29.19 $\pm$51.69 $\pm$72.36 $\pm$43.82 $\pm$35.29 $\pm$22.18 $\pm$48.52 $\pm$36.29 $\pm$34.18
[…]
$\pm$9.92 $\pm$6.91 $\pm$9.73 $\pm$6.28 $\pm$11.94 $\pm$13.42 $\pm$10.64 $\pm$6.84 $\pm$13.22 $\pm$5.91 $\pm$8.15 $\pm$5.76
[…]
$\pm$10.39 $\pm$2.61 $\pm$9.49 $\pm$11.62 $\pm$3.42 $\pm$14.74 $\pm$12.41 $\pm$3.35 $\pm$5.74 $\pm$2.45 $\pm$7.56 $\pm$6.16
[…]
$\pm$51.81 $\pm$34.83 $\pm$45.29 $\pm$80.30 $\pm$55.20 $\pm$18.91 $\pm$83.02 $\pm$40.77 $\pm$34.95 $\pm$38.79 $\pm$34.72 $\pm$26.78
[…]
$\pm$12.00 $\pm$7.47 $\pm$14.37 $\pm$17.70 $\pm$8.33 $\pm$19.11 $\pm$12.34 $\pm$3.79 $\pm$5.44 $\pm$5.92 $\pm$14.15 $\pm$6.11
[…]
$\pm$5.02 $\pm$4.60 $\pm$8.72 $\pm$15.85 $\pm$5.83 $\pm$8.05 $\pm$2.45 $\pm$4.34 $\pm$8.65 $\pm$4.77 $\pm$3.29 $\pm$5.18
[…]
$\pm$11.17 $\pm$67.14 $\pm$41.49 $\pm$91.71 $\pm$49.11 $\pm$38.97 $\pm$82.82 $\pm$63.06 $\pm$38.53 $\pm$60.10 $\pm$52.86 $\pm$76.21
[…]
$\pm$15.98 $\pm$9.66 $\pm$19.43 $\pm$43.94 $\pm$8.97 $\pm$7.17 $\pm$4.43 $\pm$9.91 $\pm$17.01 $\pm$11.93 $\pm$1.74 $\pm$18.68
Table 2: Performance comparison of adaptive teaming without teammate modeling across environments with varying difficulties. H-D denotes HOLA-Drone (V2).
@@ 4 Experiment
In this section, we assess baseline methods from our algorithm zoo to validate their practical effectiveness in multi-drone pursuit tasks across four progressively challenging environments: 4p2e3o, 4p2e1o, 4p2e5o, and 4p3e5o. The experiments are organized into two primary components: (1) adaptive teaming without explicit teammate modeling, and (2) adaptive teaming incorporating teammate modeling. Each subsection outlines the experimental setups, highlights critical findings, and provides detailed analyses.
Adaptive Teaming without Teammate Modeling. Fig. 2 and Table 2 comprehensively evaluate four baseline methods—SP, PBT, HOLA-Drone (V2) w/o g, and HOLA-Drone (V2)—across four progressively challenging multi-drone pursuit scenarios (4p2e3o, 4p2e1o, 4p2e5o, and 4p3e5o). The HOLA-Drone (V2) w/o g variant serves as an ablation study, removing the core hypergraphic game mechanism to assess its impact. Each scenario is tested against three distinct unseen teammate zoos. Fig. 2 primarily highlights success rates (SUC), while Table 2 further details collision counts (COL), average steps taken (AST), and cumulative rewards (REW). Together, these metrics demonstrate the efficacy of adaptive teaming methods within the AT-Drone benchmark.
Red dotted lines in Fig. 2 denote approximate upper-bound performances established by best-response (BR) policies specifically trained for each unseen teammate zoo, serving as critical reference points for evaluating adaptive teaming effectiveness. As environmental complexity escalates from the simplest scenario (4p2e1o) to the most complex (4p3e5o), success rates across all evaluated methods decrease, and the performance gap compared to BR policies widens. This clearly illustrates the increasing challenge of effective coordination with unfamiliar teammates.
Overall, HOLA-Drone (V2) consistently demonstrates superior performance across most adaptive teaming scenarios, achieving higher success rates, fewer collisions, and reduced average steps compared to other baseline methods. Notably, the performance advantages of HOLA-Drone (V2) become significantly more pronounced in complex scenarios, underscoring the value of advanced adaptive strategies in dynamic environments involving diverse teammate behaviors.
Interestingly, the simplified ablation model, HOLA-Drone (V2) w/o g—lacking the core hypegraphical-form game module—achieves comparable or even superior performance metrics in simpler scenario 4p2e1o. This observation suggests that while the hierarchical graphical module greatly benefits coordination in challenging environments, it may introduce unnecessary complexity when facing relatively straightforward conditions.
Adaptive teaming with teammate modelling. Fig. 3 illustrates the comparative performance of MAPPO, NAHT-D w/o Dec, and NAHT-D methods across four multi-drone pursuit scenarios: 4p2e3o, 4p2e1o, 4p2e5o, and 4p3e5o, explicitly incorporating teammate modeling into the adaptive teaming process. The NAHT-D w/o Dec variant serves as an ablation study, removing the teammate modeling decoder network and the corresponding loss function to assess their impact.
Across simpler scenarios (4p2e3o and 4p2e1o), NAHT-D achieves marginally higher or comparable success rates (SUC) compared to NAHT-D w/o Dec, with both outperforming MAPPO. However, as scenario complexity increases (4p2e5o and 4p3e5o), NAHT-D w/o Dec demonstrates notably superior performance over NAHT-D, indicating that the additional complexity introduced by the teammate modeling decoder may negatively impact coordination efficiency under highly challenging conditions. Regarding collision counts (COL) and average steps (AST), NAHT-D w/o Dec consistently achieves better or comparable results compared to NAHT-D, further suggesting that simpler teammate modeling strategies can offer superior robustness and efficiency in complex adaptive teaming scenarios. These insights are also supported by cumulative rewards (REW), where NAHT-D w/o Dec generally achieves higher or similar values across all settings.
@@
Figure 3: Performance comparison of adaptive teaming with teammate modeling across environments with varying difficulties.
Case study and demo videos. Appendix E provides a detailed case study demonstrating the real-world deployment of our adaptive teaming approach. In this case study, ATM learners are paired with unseen drone partners sampled from Zoo 3 and tasked with operating in the most challenging environment 4p3e5o. Further demonstration videos can be accessed on our project website at https://sites.google.com/view/at-drone.